In a certain federal prison, it is known that 2/3 of the inmates are under 25 years of age. It is also known that 3/5 of the inmates are male and that 6/8 of the inmates are female or 25 years of age or older. What is the probability that a prisoner selected at random from this prison is female and at least 25 years old?


1
Expert's answer
2021-01-04T20:22:50-0500

Let A be the event which denotes that the inmates are under 25 years of age.

AC denotes that the inmates are at least 25 years of age.

"P(A) = \frac{2}{3} \\

P(A^C) = 1 -P(A) \\

= 1 -\frac{2}{3} \\

= \frac{1}{3}"

Let M be the event which denotes that the inmate is male.

MC denotes that a inmate is female

"P(M) = \frac{3}{5} \\

P(M^C) = 1 -P(M) \\

= 1 -\frac{3}{5} \\

= \frac{2}{5}"

Now, we have to find the probability that a prisoner selected at random from the prison is female at least 25 years old

"P = \frac{1}{3} \times \frac{2}{5} \\

= \frac{2}{15} \\

= 0.1333"

Answer: 0.1333